You drop a rock from rest off a cliff, and it hits the ground 8.1 seconds later. How tall is the cliff?
1
Expert's answer
2019-07-24T15:48:54-0400

The distance traveled by rock

"s=v_it+\frac{gt^2}{2}"

"=0\times 8.1+\frac{9.8\times 8.1^2}{2}"

"=321\:\rm{m}"